# On complexity of restricted fragments of Decision DNNF

### v1

@@ 3.2 $\wedge_{d}$-obdd lower bounds and separations
(Theorem 5) We demonstrate that the statement of the theorem holds for the family of cnfs $\psi(G_{n},E_{hor},E_{vert})$ with $n>0$. For the lower bound, let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be a linear order obeyed by $B$. Let $\pi^{*}$ be the linear order on $V(G^{*})$ induced by $\pi_{0}$ (that is, two elements are ordered by $\pi^{*}$ in exactly the same way as they are ordered by $\pi_{0}$).
By Lemma 8, we can assume w.l.o.g, that $\pi^{*}$ is neatly crossed by a matching $M$ of $G_{n}[E_{hor}]^{*}$ of size $\Omega(lsimw(G_{n}))$. By Proposition 3, $|M|=\Omega(n)$. Then, by Theorem 9, we conclude that
@@
A $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and obeying $\pi^{*}$ is of size $2^{\Omega(n)}$.
Let ${\bf g}=\{(\textsc{jn},1)\}$ It is not hard to see that $\psi(G_{n},E_{hor},E_{vert})|_{\bf g}=\psi(G_{n}[E_{hor}])$ and the latter does not have non-essential variables. By Lemma 11, a $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and obeying $\pi^{*}$ can be obtained from $B$ by a transformation that does not increase its size By Claim 1, the size of $B$ is $2^{\Omega(n)}$ as required.
For the upper bound, we consider an fbdd whose source $rt$ is associated with jn. Let $u_{0}$ and $u_{1}$ be the children of $rt$ with respective edges $(rt,u_{0})$ and $(rt,u_{1})$ labelled with 00 and $1$. Then $u_{0}$ is the source of an fbdd representing $\psi(G_{n}[E_{vert}])$ and $u_{1}$ is the source of an fbdd representing $\psi(G_{n}[E_{hor}])$. It is not hard to see that the resulting fbdd indeed represents $\psi(G_{n},E_{hor},E_{vert})$. It remains to demonstrate that both $\psi(G_{n}[E_{hor}])$ and $\psi(G_{n}[E_{vert}])$ can be represented by poly-size fbdds. We demonstrate that the incidence pathwidth of both $\psi(G[E_{hor}])$ and $\psi(G_{n}[E_{hor}])$ and $\psi(G_{n}[E_{vert}])$ is at most $7$. It then follows [13] that both these cnfs can be represented by linear-size fbdds.
@@
If $u=(i,1)$ for $1\leq i\leq n$ then ${\bf B}(u)=\{u[1],u[2],\neg V(G)[1],\neg V(G)[2]\}$ (recall that the last two elements correspond to the long clauses). If $u=(i,j)$ for $1\leq i\leq n$ and $2\leq j\leq n$, let $v=(i,j-1)$. Then ${\bf B}(u)=\{u[1],u[2],v[1],v[2],(u[1]\vee v[2]),(u[2]\vee v[1]),\neg V(G)[1],\neg V(G)[2]\}$. It follows from a direct inspection that $(P,{\bf B})$ is a tree decomposition of the incidence graph of $\psi(G_{n}[E_{hor}])$ of width $7$. For $\psi(G_{n}[E_{vert}])$, the construction is analogous with rows and columns changing their roles. ∎
@@
(Theorem 6) We consider the class of cnfs $\varphi(G_{n},E_{hor},E_{vert})$. Let $B$ be an obdd representing $\varphi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be an order obeyed by $B$ and let $\pi$ be a linear order of $V(G_{n})$ induced by $\pi_{0}$. By Lemma 8, we can assume w.l.o.g. the existence of a matching $M$ of $\varphi(G_{n}[E_{1}])$ of size at least $\Omega(n)$ crossing $\pi$. By Theorem 10, an obdd obeying $\pi$ and representing $\varphi(G_{n}[E_{1}])$ is of size $2^{\Omega(n)}$. As all the variables of $\varphi(G_{n}[E_{1}])$ are essential, It follows from Lemma 11 that an obdd representing $\varphi(G_{n}[E_{1}])$ can be obtained from $B$ by a transformation that does not increase its size (note that the transformation does not introduce conjunction nodes hence, being applied to an obdd produces another obdd). We thus conclude that the size of $B$ is at least $2^{\Omega(n)}$.
For the upper bound, we demonstrate existence of a poly-size $\wedge_{d}$-obdd representing $\varphi(G_{n}E_{hor},E_{vert})$ and obeying the order $\pi_{d}$ where the first element is jn followed by $V(G_{n})$ in the dictionary order, that is $(i_{1},j_{1})$ precedes $(i_{2},j_{2})$ if $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$.
@@
There are linear sized obdds obeying $\pi_{d}$ and representing $\varphi(Hor_{i})$ and $\varphi(Vert_{i})$ for each $1\leq i\leq n$.
@@
In order to represent $\mathsf{var}(Hor_{i})$, let us query the variables by the ascending order of their second coordinate, that is $(i_{1},),\dots,(i,n)$. For each $11\leq j\leq n-$ 1, let ${\bf H}_{j}$ be the set of all functions $\varphi(Hr_{i})|_{\bf g}$ where ${\bf g}$ is an assignment over $(i,1),\dots,(i,j)$. It is not hard to see that that there are at most $3$ such functions: constant zero, and two functions on $(i,j+1),\dots,(i,n)$ completely determined by the assignment of $(i,j)$. It is well known [20] that the resulting obdd can be seen as a dag with layers $11,\dots,n+$ 1, the last layer is reserved for the sink nodes, and each layer $1\leq j\leq n$ containing nodes labelled with $(i,j)$ and having $O(1)$ size.
@@
Let $B^{h}_{i}$ and $B^{v}_{i}$ be obdds obeying $\pi_{d}$ and representing $Hor_{i}$ and $Vert_{i}$ as per Claim 2. By using $O(n)$ conjunction nodes, it is easy to create $\wedge_{d}$-obdds $B_{1}$ and $B_{0}$ that are, respectively conjunctions of all $B^{v}_{i}$ and all $B^{h}_{i}$ and hence, respectively represent $\wedge_{1\leq i\leq n}\varphi(Hor_{i})$ and $\wedge_{1\leq i\leq n}\varphi(Vert_{i})$. It is not hard to see that the former conjunction is $\varphi(G[E_{hor}])$ and the latter one is $\varphi(G[E_{vert}])$, hence they are, respectively represented by $B_{1}$ and $B_{0}$. The resulting $\wedge_{d}$-obdd is then constructed as follows. The source $rt$ is a decision node labelled with jn. Let $u_{1}$ and $u_{0}$ be the children of $rt$ labelled by $1$ and 00 respectively. Then $u_{1}$ is the source of $B_{1}$ and $u_{0}$ is the source of $B_{0}$. ∎
@@
Note that the variables of a single column of $G_{n}$ explored along the column satisfy the dictionary order. However, if we try to explore first one column and then another one then the dictionary order will be violated. Thus, the upper bound in the proof of Theorem 6 demonstrates how the conjunction nodes overcome the rigidity of obdd by splitting the set of all variables in chunks that do obey a fixed order.
3.3 Proofs of Theorem 9 and 10
We first prove Theorem 9 using the approach as outlined in Section 3.1 . The proof of Theorem 10 will follow the same approach but with a simplified reasoning.
The first step of the approach is to define a fooling set of assignments. W.l.o.g. we assume existence of sets $U,W\subseteq V(G)$. such that $M$ is a matching between $U[1]$ and $W[2]$ and that each element of $U[1]$ is ordered by $\pi^{*}$ before each element of $W[2]$. We let $\pi_{0}$ be the prefix of $\pi^{*}$ whose last element is the last element of $U[1]$ in $\pi^{*}$.
@@ 3.3 Proofs of Theorem 9 and 10
An assignment ${\bf g}$ over $\pi_{0}$ is fooling if it satisfies the following two conditions.
@@
    At least one variable of $U[1]$ is assigned with 00 and at least two variables of $U[1]$ are assigned with $1$.
@@
Before we proceed, we extend the notation by letting $U=\{u_{1},\dots,u_{q}\}$, $W=\{w_{1},\dots,w_{q}\}$ and $M=\{\{u_{1}[1],w_{1}[2]\},\dots,\{u_{q}[1],w_{q}[2]\}\}$. Further on, for $I\subseteq\{1,\dots,q\}$, we let $U_{I}=\{u_{i}|i\in I\}$ and $U[1]_{I}=\{u_{i}[1]|i\in I\}$. The sets $W_{I}$ and $W[2]_{I}$ are defined accordingly.
We now proceed to the next stage, that is, defining unbreakable sets.
@@
Let ${\bf g}\in\mathcal{F}$. We denote by $I({\bf g})$ the set of all $1\leq i\leq q$ such that $u_{i}[1]$ is assigned with $1$ by ${\bf g}$. We call $W[2]_{I({\bf g})}$ the unbreakable set of ${\bf g}$. Further on for $J\subseteq I({\bf g})$, we let ${\bf h}_{J}[{\bf g}]$ to be the assignment over $V(G^{*})$ such that ${\bf g}\subseteq{\bf h}_{J}[{\bf g}]$, for each $u\in V(G^{*})$, ${\bf h}_{J}[{\bf g}]=0$ if $u\in W[2]_{J}$ and $1$ otherwise.
Note that we introduce a notation for unbrekable sets specifically tailored to demonstrate the role of $I({\bf g})$.
@@
Continuing on Example 4, let ${\bf g}=\{(u_{1}[1],0),(u_{2}[1],1),(u_{3}[1],1),\\ (u_{1}[2],1),(u_{2}[2],1),(w_{1}[1],1),(w_{2}[1],1)\}$. Then $I({\bf g})=\{2,3\}$ and the unbreakable set for ${\bf g}$ is $\{w_{2}[2],w_{3}[2]\}$. For $J=I({\bf g})$, ${\bf h}_{J}({\bf g})$ is the extension of ${\bf g}$ assigning $w_{1}[2]$ with $1$, $w_{2}[2]$ with 00 and $w_{3}[2]$ with 00.
The next stage of the proof is to show that the unbreakable sets are indeed unbrekable by $\mathcal{S}(B)|_{\bf g}$. We need first an auxiliary statement.
@@
${\bf h}_{J}[{\bf g}]$ satisfies $\psi(G)$ if and only if $J\neq\emptyset$.
@@
First of all, if $J=\emptyset$ then $U[2]$ is assigned with $1$ for each $u\in V(G)$ thus falsifying the clause $\neg V(G)[2]$. Otherwise, we observe that $\neg V(G)[1]$ is satisfied because one of $U[1]$ is assigned negatively and $\neg V(G)[2]$ is satisfied because a variable of $W[2]_{J}$ is assigned negatively. Hence, $\psi(G)$ may be falsified only if both variables of a cause $(u[1]\vee w[2])$ are assigned with zeroes. By assumption $w[2]=w_{j}[2]$ for some $j\in J$. Further on, $u_{j}[1]$ assigned with $1$ by ${\bf g}$, but on the other hand, the only variables assigned by zeroes belong to $U[1]$. Hence $u[1]=u_{i}[1]$ for some $i\neq j$. By definition of $\psi(G)$, $G^{*}$ has an edge $\{u_{i}[1],w_{j}[2]\}$ in contradiction to $M$ being an induced matching. ∎
Now, we are ready to prove the unbreakbility of the unbreakable sets.
@@
For each ${\bf g}\in\mathcal{F}$, $\mathcal{S}(B)|_{\bf g}$ does not break $W[2]_{I({\bf g})}$.
@@
For the existence $u({\bf g})$, we now need to rule out the possibility that $W[2]_{I({\bf g})}\subseteq X$ where $X$ is as in Lemma 3.
@@
Let ${\bf g}\in\mathcal{F}$. Then there is $u\in L({\bf g})$ such that $W[2]_{I({\bf g})}\subseteq\mathsf{var}(B_{u})$.
@@
It follows from the combination of Lemma 3 that either the statement of the present lemma is true or $W[2]_{I({\bf g})}\subseteq X$, where $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{u\in L_{\bf g}}\mathsf{var}B_{u}$.
First all, we observe that by Lemma 12 $\mathcal{S}(B)|_{\bf g}$ includes at least one satisfying assignment, namely ${\bf h}_{J}[{\bf g}]$ where $J$ is an arbitrary non-empty subset of $I({\bf g})$. Next, we observe that $X$ cannot include all the variables of $\mathcal{S}(B)|_{\bf g}$ as otherwise, assignment all of them with $1$ will falsify $\neg V(G)[2]$. It then follows from Lemma 3 that $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\mathcal{S}$ where $\mathsf{var}(\mathcal{S})\neq\emptyset$.
By definition, $|W[2]_{I({\bf g})}|\geq 2$. Therefore, if $W[2]_{I({\bf g})}\subseteq X$, then we can partition $X$ into $X_{1}$ and $X_{2}$ each $X_{i}$ has a non-empty intersection with $|W[2]_{I({\bf g})}|$. But then $\mathcal{S}(B)|_{\bf g}=X_{1}^{\{0,1\}}\times X_{2}^{\{0,1\}}\times\mathcal{S}$ thus breaking $W[2]_{I({\bf g})}$ in contradiction to Lemma 13. ∎
In light of Lemma 14, we now define $u({\bf g})$ as the $u\in L({\bf g})$ such that $W[2]_{I({\bf g})}\subseteq\mathsf{var}(B_{u})$.

### v2

@@ 3.2 $\wedge_{d}$-OBDD lower bounds and separations
(Theorem 6) We demonstrate that the statement of the theorem holds for the family of cnfs $\psi(G_{n},E_{hor},E_{vert})$ with $n>0$. For the lower bound, let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be a linear order respected by $B$. Let $\pi^{*}$ be the linear order on $V(G^{*})$ induced by $\pi_{0}$ (that is, two elements are ordered by $\pi^{*}$ in exactly the same way as they are ordered by $\pi_{0}$).
By Lemma 9, we can assume w.l.o.g, that $\pi^{*}$ is neatly crossed by an induced matching $M$ of $G_{n}[E_{hor}]^{*}$ of size $\Omega(lsimw(G_{n}))$. By Proposition 3, $|M|=\Omega(n)$. Then, by Theorem 10, we conclude that
@@
A $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and respecting $\pi^{*}$ is of size $2^{\Omega(n)}$.
Let ${\bf g}=\{(\textsc{jn},1)\}$ It is not hard to see that $\psi(G_{n},E_{hor},E_{vert})|_{\bf g}=\psi(G_{n}[E_{hor}])$ and the latter does not have non-essential variables. By Lemma 12, a $\wedge_{d}$-obdd representing $\psi(G_{n}[E_{hor}])$ and respecting $\pi^{*}$ can be obtained from $B$ by a transformation that does not increase its size By Claim 1, the size of $B$ is $2^{\Omega(n)}$ as required.
For the upper bound, we consider an fbdd whose source $rt$ is associated with jn. Let $u_{0}$ and $u_{1}$ be the children of $rt$ with respective edges $(rt,u_{0})$ and $(rt,u_{1})$ labelled with 0 and $1$. Then $u_{0}$ is the source of an fbdd representing $\psi(G_{n}[E_{vert}])$ and $u_{1}$ is the source of an fbdd representing $\psi(G_{n}[E_{hor}])$. It is not hard to see that the resulting fbdd indeed represents $\psi(G_{n},E_{hor},E_{vert})$. It remains to demonstrate that both $\psi(G[E_{hor}])$ $\psi(G_{n}[E_{vert}])$ can be represented by poly-size fbdds. We demonstrate that the incidence pathwidth of both $\psi(G_{n}[E_{hor}])$ and $\psi(G_{n}[E_{vert}])$ is at most $7$. It then follows [17] that both these cnfs can be represented by linear-size fbdds.
@@
If $u=(i,1)$ for $1\leq i\leq n$ then ${\bf B}(u)=\{u[1],u[2],\neg V(G)[1],\neg V(G)[2]\}$ (recall that the last two elements correspond to the long clauses). If $u=(i,j)$ for $1\leq i\leq n$ and $2\leq j\leq n$, let v=(i,j1)v=(i,j-1). Then ${\bf B}(u)=\{u[1],u[2],v[1],v[2],(u[1]\vee v[2]),(u[2]\vee v[1]),\neg V(G)[1],\neg V(G)[2]\}$. It follows from a direct inspection that $(P,{\bf B})$ is a tree decomposition of the incidence graph of $\psi(G_{n}[E_{hor}])$ of width $7$. For $\psi(G_{n}[E_{vert}])$, the construction is analogous with rows and columns changing their roles. ∎
@@
(Theorem 7) We consider the class of cnfs $\varphi(G_{n},E_{hor},E_{vert})$. Let $B$ be an obdd representing $\varphi(G_{n},E_{hor},E_{vert})$. Let $\pi_{0}$ be an order respected by $B$ and let $\pi$ be a linear order of $V(G_{n})$ induced by $\pi_{0}$. By Lemma 9, we can assume w.l.o.g. the existence of an induced matching $M$ of $\varphi(G_{n}[E_{hor}])$ of size at least $\Omega(n)$ crossing $\pi$. By Theorem 11, an obdd respecting $\pi$ and representing $\varphi(G_{n}[E_{hor}])$ is of size $2^{\Omega(n)}$. As all the variables of $\varphi(G_{n}[E_{hor}])$ are essential, it follows from Lemma 12 that an obdd representing $\varphi(G_{n}[E_{hor}])$ can be obtained from $B$ by a transformation that does not increase its size (note that the transformation does not introduce conjunction nodes hence, being applied to an obdd produces another obdd). We thus conclude that the size of $B$ is at least $2^{\Omega(n)}$.
For the upper bound, we demonstrate existence of a poly-size $\wedge_{d}$-obdd representing $\varphi(G_{n}E_{hor},E_{vert})$ and respecting the order $\pi_{d}$ where the first element is jn followed by $V(G_{n})$ in the dictionary order, that is $(i_{1},j_{1})$ precedes $(i_{2},j_{2})$ if $i_{1}<i_{2}$ or $i_{1}=i_{2}$ and $j_{1}<j_{2}$.
@@
There are linear sized obdds respecting $\pi_{d}$ and representing $\varphi(Hor_{i})$ and $\varphi(Vert_{i})$ for each $1\leq i\leq n$.
@@
In order to represent $\mathsf{var}(Hor_{i})$, let us query the variables by the ascending order of their second coordinate, that is $(i,1),\dots,(i,n)$. For each $1\leq j\leq n-1$, let ${\bf H}_{j}$ be the set of all functions $\varphi(Hor_{i})|_{\bf g}$ where $\bf g$} is an assignment over $(i,1),\dots,(i,j)$. It is not hard to see that that there are at most $3$ such functions: constant zero, and two functions on $(i,j+1),\dots,(i,n)$ completely determined by the assignment of $(i,j)$. It is well known [25] that the resulting obdd can be seen as a dag with layers $1,\dots,n+1$, the last layer is reserved for the sink nodes, and each layer $1\leq j\leq n$ containing nodes labelled with $(i,j)$ and having $O(1)$ size.
@@
Let $B^{h}_{i}$ and $B^{v}_{i}$ be obdds respecting $\pi_{d}$ and representing $Hor_{i}$ and $Vert_{i}$ as per Claim 2. By using $O(n)$ conjunction nodes, it is easy to create $\wedge_{d}$-obdds $B_{1}$ and $B_{0}$ that are, respectively conjunctions of all $B^{v}_{i}$ and all $B^{h}_{i}$ and hence, respectively represent $\wedge_{1\leq i\leq n}\varphi(Hor_{i})$ and $\wedge_{1\leq i\leq n}\varphi(Vert_{i})$. It is not hard to see that the former conjunction is $\varphi(G[E_{hor}])$ and the latter one is $\varphi(G[E_{vert}])$, hence they are, respectively represented by $B_{1}$ and $B_{0}$. The resulting $\wedge_{d}$-obdd is then constructed as follows. The source $rt$ is a decision node labelled with jn. Let $u_{1}$ and $u_{0}$ be the children of $rt$ labelled by $1$ and 0 respectively. Then $u_{1}$ is the source of $B_{1}$ and $u_{0}$ is the source of $B_{0}$. ∎
@@
Note that the variables of a single column of $G_{n}$ explored along the column satisfy the dictionary order. However, if we try to explore first one column and then another one then the dictionary order will be violated. Thus, the upper bound in the proof of Theorem 7 demonstrates how the conjunction nodes overcome the rigidity of obdd by splitting the set of all variables in chunks that do respect a fixed order.
3.3 Proofs of Theorem 10 and 11
In this subsection we prove Theorems 10 and 11 using the approach as presented in Section 3.1.
We first prove Theorems 10. For that purpose, we define a set $\mathcal{F}$ of assignments of $\psi(G)$ along with their $UB$ sets. Then we demonstrate that $\mathcal{F}$ is a fooling set and its size is exponential in $|M|$. The required lower bound will then immediately follow from Lemma 4. W.l.o.g. we assume existence of sets $U,W\subseteq V(G)$. such that $M$ is a matching between $U[1]$ and $W[2]$ and that each element of $U[1]$ is ordered by $\pi^{*}$ before each element of $W[2]$. We let $\pi_{0}$ be the prefix of $\pi^{*}$ whose last element is the last element of $U[1]$ in $\pi^{*}$.
@@ 3.3 Proofs of Theorem 10 and 11
We define $\mathcal{F}$ as the set of all assignments over $\pi_{0}$ satisfying the following two conditions.
@@
    At least one variable of $U[1]$ is assigned with 0 and at least one variable of $U[1]$ is assigned with $1$.
@@
Before we proceed, we extend the notation by letting $U=\{u_{1},\dots,u_{q}\}$, $W=\{w_{1},\dots,w_{q}\}$ and $M=\{\{u_{1}[1],w_{1}[2]\},\dots,\{u_{q}[1],w_{q}[2]\}\}$. Further on, for $I\subseteq\{1,\dots,q\}$, we let $U_{I}=\{u_{i}|i\in I\}$, $U[1]_{I}=\{u_{i}[1]|i\in I\}$, and $U[2]_{I}=\{u_{i}[2]|i\in I\}$. The sets $W_{I}$, $W[1]_{I}$, and $W[2]_{I}$ are defined accordingly.
@@
Let $\bf g$$\in\mathcal{F}$. We denote by $I({\bf g})$ the set of all $1\leq i\leq q$ such that $u_{i}[1]$ is assigned with $1$ by $\bf g$}. We refer to $W[2]_{I({\bf g})}$ as $UB({\bf g})$.
We are now going to prove that the assignments $\bf g$$\in\mathcal{F}$ equipped with the sets $UB({\bf g})$ form a fooling set. For this purpose, we will use assignments of a special form as defined below.
Definition 17.
Let $\bf g$$\in\mathcal{F}$. Let $J\subseteq I({\bf g})$. We denote by ${\bf h}_{J}[{\bf g}]$ the assignment over $V(G^{*})$ such that ${\bf g}\subseteq{\bf h}_{J}[{\bf g}]$, for each $u\in V(G^{*})\setminus\mathsf{var}({\bf g})$, ${\bf h}_{J}[{\bf g}]=0$ if $u\in W[2]_{J}$ and 11 otherwise.
@@
Continuing on Example 6, let ${\bf g}=\{(u_{1}[1],0),(u_{2}[1],1),(u_{3}[1],1),\\ (u_{1}[2],1),(u_{2}[2],1),(w_{1}[1],1),(w_{2}[1],1)\}$. Then $I({\bf g})=\{2,3\}$ and $UB({\bf g})$ is $\{w_{2}[2],w_{3}[2]\}$. For $J=I({\bf g})$, ${\bf h}_{J}({\bf g})$ is the extension of $\bf g$} assigning $w_{2}[2]$ and $w_{3}[2]$ with 0 and assigning with $1$ the remaining variables that are not assigned by $\bf g$}.
We are now proving an auxiliary lemma about the assignments ${\bf h}_{J}[{\bf g}]$.
@@
Let $\bf g$$\in\mathcal{F}$. Let $J\subseteq I({\bf g})$. ${\bf h}_{J}[{\bf g}]$ satisfies $\psi(G)$ if and only if $J\neq\emptyset$. In particular, ${\bf h}_{I}[{\bf g}]$ satisfies $\psi(G)$.
@@
First of all, if $J=\emptyset$ then ${\bf h}_{J}[{\bf g}]$ falsifies the clause $\neg V(G)[2]$. Indeed, the only variables of $V(G)[2]$ assigned by ${\bf h}_{J}[{\bf g}]$ with 0 are those of $W[2]_{J}$. But $J=\emptyset$ implies that $W[2]_{J}=\emptyset$.
Otherwise, we observe that $\neg V(G)[1]$ is satisfied because one of $U[1]$ is assigned negatively and $\neg V(G)[2]$ is satisfied because a variable of $W[2]_{J}$ is assigned negatively. Hence, $\psi(G)$ may be falsified only if both variables of a clause $(u[1]\vee w[2])$ are assigned with zeroes. By assumption $w[2]=w_{j}[2]$ for some $j\in J$. Further on, $u_{j}[1]$ assigned with $1$ by $\bf g$}, but on the other hand, the only variables of $V(G)[1]$ assigned by zeroes belong to $U[1]$. Hence $u[1]=u_{i}[1]$ for some $i\neq j$. By definition of $\psi(G)$, $G^{*}$ has an edge $\{u_{i}[1],w_{j}[2]\}$ in contradiction to $M$ being an induced matching.
For the second statement, we note that $I({\bf g})\neq\emptyset$ because, by definition of $\bf g$}, at least one variable of $U[1]$ is assigned with $1$. Then we apply the first statement, ∎
Now we are ready to prove that the $UB$ sets are indeed unbreakable.
@@
For each $\bf g$$\in\mathcal{F}$, $\mathcal{S}(B)|_{\bf g}$ does not break $UB({\bf g})=W[2]_{I({\bf g})}$.
@@
We next prove that the distinct projections property holds for $\mathcal{F}$.
@@
Let ${\bf g}_{1},{\bf g}_{2}$ be two distinct elements of $\mathcal{F}$. Then $Proj(\mathcal{S}(B)|_{{\bf g}_{1}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))\neq Proj(\mathcal{S}(B)|_{{\bf g}_{2}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$.